Donald derives utility from only two goods, carrots (X) and donuts (Y).

 His utility function is as follows: U(X,Y) = X1/4Y3/4. Donald has an income (M) of $120 and the price of carrots (PX) is $2 while the price of donuts (PY) is $6. What quantities of carrots and donuts will maximize Donald's utility? How does MRSXY change as the firm uses more Xholding utility constant




1
Expert's answer
2021-04-18T19:29:42-0400

SOLUTION.

Quantity of Donuts and Carrots that maximizes Donald’s utility.


Utility Function= "U" "(X,Y)" ="X^{1\/4}\nY^{3\/4}"


Income "(M)" = $120.


Price of Carrots "P_{x}" = $2

Price of Donuts "P_{y}" = $6


The marginal rate of substitution ("MRS" ) is the rate, which the consumer is willing to substitute one good for another.

In this case, "MRS" = "\frac{P_x}{P_y}"

"\frac{P_x}{P_y}"  is the price ratio of carrots and donuts.


The budget line will be,

Income ("M" ) "=" (price of carrots "\times" quantity of carrots) "+" (price of donuts "\times" quantity of donuts)

"M" "=" ("P_x" "\times Q_y" ) "+" ("P_y \times Q_y)"

"120" "=" "( 2 \times Q_x) + (6 \times Q_y)"

"120 = 2Q_x + 6Q_y"


"MRS" "= \frac{marginal utility of carrots}{marginal utility of donuts}"


"MRS = \frac{MU_x}{MU_y}"


"U = X^\frac{1}{4}Y^\frac{3}{4}"


"MU_x = \frac{du}{dx} = \frac{1}{4}X^\frac{-3}{4}Y^\frac{3}{4}"


"MU_y = \frac{du}{dy} = \frac{3}{4}X^\frac{1}{4}Y^\frac{-1}{4}"


"MRS = (\frac{1}{4}X^\frac{-3}{4}Y) / ( \frac{3}{4}X^\frac{1}{4}Y^\frac{-1}{4} )" simplify the equation


"MRS = \frac{Y}{3X}"

"MRS = \frac{Y}{3X} = \frac{P_y}{P_x} = \frac{6}{2}"  


          "Y=9X"


Substitute "Y=9X" into the Budget line equation

"120=2Q_x+6Q_y"

"120=2Q_x+6Q(9x)"

"120=2Q_x+54Q_x"

"120=56Q_x"

"Q_x=\frac{15}{7}"


Substitute "Q_x" to the equation to find "Q_y"

"120=2(\frac{15}{7})+6Q_y"

"120= \frac{30}{7}+ 6Q_y"

"Q_y=\frac{135}{7}"


Therefore the Quantities that will maximize Donald’s utility are "Q_x=\frac{15}{7}" and "Q_y=\frac{135}{7}"


How does MRSxy change as the firm uses more X, holding utility constant?

  • Since the firm will use more X and less Y, the "MRS>\frac{P_x}{P_y}"

Therefore, MRS will be greater.